OpenStudy (anonymous):

How do I invert this matrix (or show that it is singular)? (1 4 -2 3)

OpenStudy (anonymous):

If you mean inverse by invert then: if the Determinant=0 then it doesn't have an inverse So let: \[A=\left[\begin{matrix}a & b \\ c & d\end{matrix}\right] \quad \text{then} \quad A^-1=\frac{ 1 }{ detA }\left[\begin{matrix}d & -b \\ -c & a\end{matrix}\right]\] if you don't then sorry! :-)

OpenStudy (anonymous):

det(a)=ad-bc

OpenStudy (anonymous):

The determinant is 3-(-8) = 11. So i scalar multiply the matrix (1 -4 2 3) by 1/11?

OpenStudy (anonymous):

yes, you just need to swap the 1 and 3 around (check, a and d swap!) but yep that's the right idea

OpenStudy (anonymous):

Woops, I meant (3 -4 2 1) Yes, so I will get a lot of decimals?

OpenStudy (anonymous):

just leave it in fraction form if I was you, more accurate technically

OpenStudy (anonymous):

Ok... one more thing, what would be the case if the matrix was singular? is that when the determinant is zero?

OpenStudy (anonymous):

Exactly!